DIRC Operations Weekly Report

Period: Saturday 2006/05/27 - Friday 2006/06/02

Oncall: Emmanuel Latour, Jerome Bequilleux
Fast Monitoring: Emmanuel Latour, Jerome Bequilleux

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060602.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060602/ (empty)

Contents


BaBar-PEP news

PEP peak lumi: 10.4 1033 (previous week: 10.87 1033)

BaBar integrated lumi: 2.669 fb-1 (previous week: 1.389 fb-1)

Run 5b PEP delivery: 52.3 fb-1

Run 5 total delivery: 110.74 fb-1 Run 5 total BaBar recorded: 106.54 fb-1 including 5.55 fb-1 off-peek data

PEP activities this week:

  • LER characterised
  • Problem with QD4R power supply on going (started the 25th of May)
  • Problems in the damping ring+ on the magnets that enable to inject electrons in the ring

    PEP activities planned next week :

  • characterization on Tuesday morning
  • Off peak data taking for a week just after the characterization

    Foreseen work opportunities (PEP down periods): None.

    BaBar activities/issues this week: nothing special

    DIRC activities

    See DIRC logbook for details

    Downtime due to the DIRC: None.

    Backgrounds

    Background summary plot

    Changed the scalers thresholds due to high backgrounds with the high luminosity PEP has reached.
    Previous values:

    New values:

    Scheduled activities

    Pending Work

    Links